In Capsim, investment in plant & equipment (including automation) is depreciated on a straight line basis over 15 years. Imagine the firm you inherited has 1st shift capacity of 1,200,000 units at an automation level of 5. The original cost of this capacity was $31.2 million. Because the machines are 4 years old, 4/15 of the value has been depreciated. Because you only currently sell 800,000 units of the related product per year, you decide to sell 600,000 units of capacity. Selling capacity recoups 65% of the original purchase price. What will be the gain/loss for this transaction on the income statement?

a) $10,140,000

b) $4,160,000

c) -$1,300,000

d) $15,600,000
Answer #1

Original Cost = 31200000

WDV after 4 years = Original Cost*11/15 = 31200000*11/15 = 22880000

WDV for 600000 capacity = Total WDV*600000/Total Capacity = 22880000*600000/1200000 = $11440000

Selling Price = 65% of Original Cost(for 600000 capacity) = 65% of 31200000*600000/1200000 = $10140000

Profit/(Loss) = Selling Price - WDV = 10140000-11440000 = -$1300000